OpenStudy (anonymous):

The distribution function of random variable X is given by F(X) = 1 - (1+X) e^-X >=0 then how to find density function, mean and variance of X ?

hartnn (hartnn):

@prathameshtambe007 Hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) Do you know that f(x) = F'(x) ? where f(x) is density function and F(x) is distribution function. so, you just need derivative of F(x), can you find ?

OpenStudy (anonymous):

okay thats great I'll differentiate it. but then what about mean and variance?

hartnn (hartnn):

mean is dsfined as \(\huge m=\int_D xf(x)dx\) where D is the domain of the function f(x)

hartnn (hartnn):

and variance is \(\huge V =\int_D x^2f(x)dx-m^2\)

OpenStudy (anonymous):

yep this one I know but what should be the domain? is it from 0 to infinity? also what is f(x) a distribution function or a density function? I tried this thing with integrating distribution function with from 0 to infinity and I got the answer infinity which is not acceptable. What should I do?

hartnn (hartnn):

f(x) = density function. domain = 0 to infinity

hartnn (hartnn):

what you got after differentiating ?

OpenStudy (anonymous):

oh so its density function and I tried with distribution function. Just a moment I'll try is again and let you know

hartnn (hartnn):

because i am getting a constant answer for mean.

hartnn (hartnn):

take your time :)

OpenStudy (anonymous):

I got X.e^-X as density function. is it correct?

hartnn (hartnn):

absolutely! :)

OpenStudy (anonymous):

oops then after multiplying it by X and integrating m= integration 0 to infinity X^2. e^-X = X^2 (- e^-X) - 2X (e^-X) + 2(-e^ -X) limits 0 to infinity so if I put infinity at X I'll get 0 and if I punt 0 still I'll get 0 so the answer is zero. is it so

hartnn (hartnn):

'if i put 0' <------try again everything else is correct ^_^

OpenStudy (anonymous):

oh yes got it the last term will be -2

hartnn (hartnn):

so, m = 0- (-2) = 2 is the mean.

hartnn (hartnn):

try variance now :)

OpenStudy (anonymous):

variance is 6 ???

hartnn (hartnn):

just one more look at the formula, what you got as 6 ? and then, what will be the variance ?

OpenStudy (anonymous):

shot forgot to substract m2 . varience will 6-4 = 2

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

hartnn (hartnn):

good work!

OpenStudy (anonymous):

Thanks a lot........ !!!!! @hartnn

hartnn (hartnn):

welcome ....!!! @prathameshtambe007 oh, and since you are new here, if you have any doubts browsing this site, you can ask me, or to an ambassador (with an encircled A before their name). Have fun learning with Open Study! :) Hope you like this site.

OpenStudy (anonymous):

(Y)

OpenStudy (anonymous):

thanks